# Debiased Machine Learning and Network Cohesion for Doubly-Robust Differential Reward Models in Contextual Bandits

### v2

@@ G.2 Preliminaries
G.3 Proof of Theorem 1
The proof follows closely from Abeille & Lazaric (2017) with several adjustments. Assumption 2 implies that we only need to consider the unit ball $\mathcal{X}=\{\|x\|\leq 1\}$. Then the regret can be decomposed into
$\underbrace{\sum_{k=1}^{K}\frac{1}{k}\sum_{(i,t)\in\mathcal{O}_{k}\backslash\mathcal{O}_{k-1}}\left((\phi(x_{i,t}^{\star}))^{\top}\Theta^{\star}_{k}-\phi(x_{i,t})^{\top}\tilde{\Theta}_{k}\right)}_{R^{TS}(K)}+\underbrace{\sum_{k=1}^{K}\frac{1}{k}\sum_{(i,t)\in\mathcal{O}_{k}\backslash\mathcal{O}_{k-1}}\left(\phi(x_{i,t})^{\top}\tilde{\Theta}_{k}-\phi(x_{i,t})^{\top}\Theta^{\star}_{k}\right)}_{R^{RLS}(K)}$
where $\phi(x_{i,t}^{\star})$ is the context vector under the optimal action and $\Theta^{\star}_{k}$ is the true parameter value. The first term is the regret due to the random deviations caused by sampling $\tilde{\Theta}_{k}$ and whether it provides sufficient useful information about the true parameter $\Theta^{\star}_{k}$. The second term is the concentration of the sampled term around the true linear model for the advantage function.
@@ G.3 Proof of Theorem 1
The first term describes the deviation of the TS linear predictor from the RLS one, while the second term describes the deviation of the RLS linear predictor from the true linear predictor. The first term is controlled by the construction of the sampling distribution $D^{TS}$, while the second term is controlled by the RLS estimate being a minimizer of the regularized cumulative squared error in (5). In particular, the first term will be small when the TS estimate concentrates around the RLS one, while the second will be small when the RLS estimate concentrates around the true parameter vector. The next proposition gives a lower bound on the probability that, for all stages, both the RLS parameter vector concentrates around the true parameter vector and the TS parameter vector concentrates around the RLS one.
@@
$P(\hat{E}_{K})$ $=P(\cap_{k=1}^{K}\{\|\Theta_{k}-\Theta^{\star}\|_{V_{k}}\leq\beta_{k}(\delta^{\prime})\})$
$=1-\sum_{k=1}^{K}P(\|\Theta_{k}-\Theta^{\star}\|_{V_{k}}>\beta_{k}(\delta^{\prime}))$
@@
$+\sum_{k=1}^{K}\frac{1[E_{K}]}{k}\left[\sum_{(i,t)\in\mathcal{O}_{k}\backslash\mathcal{O}_{k-1}}|\phi(x_{i,t})^{\top}(\hat{\Theta}_{k}-\Theta_{k}^{*})|\right]$
@@
$\leq\left(\beta_{K}(\delta^{\prime})+\gamma_{K}(\delta^{\prime})\right)\sqrt{H_{K}}\sqrt{2c\left[2Kd\log\left(\gamma+\lambda M+\frac{K+1}{8d}\right)-\log\det\left(V_{0}\right)\right]}$
@@
Leveraging Abeille & Lazaric (2017), Definition 1 lets us bound $R^{TS}(K)$ under the event $E_{k}$ by
@@
$R^{TS}(K)\leq\frac{4\gamma_{K}(\delta^{\prime})}{d}\left(\sqrt{\frac{8K}{\lambda}\log\left(\frac{4}{\delta}\right)}+\sqrt{4cH_{K}Kd\log\left(\gamma+\lambda M+\frac{K+1}{8d}\right)-\log\det\left(V_{0}\right)}\right).$
@@
$\begin{split}&\left(\beta_{K}(\delta^{\prime})+\gamma_{K}(\delta^{\prime})\left[1+\frac{4}{d}\right]\right)\sqrt{4cH_{K}Kd\log\left(\gamma+\lambda M+\frac{K+1}{8d}\right)-\log\det\left(V_{0}\right)}\\ +&\frac{4\gamma_{K}(\delta^{\prime})}{p}\sqrt{\frac{8K}{\lambda}\log\left(\frac{4}{\delta}\right)}\end{split}$

### v3

@@ F.2 Preliminaries
F.3 Proof of Theorem 1
We first decompose the regret bound
$\sum_{k=1}^{K}\frac{1}{k}\sum_{(i,t)\in\mathcal{O}_{k}\backslash\mathcal{O}_{k-1}}\big{[}$ $\pi_{i,t}^{\star}x(S_{i,t},a_{i,t}^{*})^{\top}\theta^{*}_{i,t}-\pi_{i,t}x(S_{i,t},A_{i,t})^{\top}\theta^{*}_{i,t}\big{]}$
$=\sum_{k=1}^{K}\frac{1}{k}\sum_{(i,t)\in\mathcal{O}_{k}\backslash\mathcal{O}_{k-1}}\big{[}$ $(\pi_{i,t}^{\star}-\pi_{i,t})x(S_{i,t},A_{i,t})^{\top}\theta^{*}_{i,t}$
$+\pi_{i,t}^{\star}\left\{x(S_{i,t},a_{i,t}^{*})^{\top}\theta^{*}_{i,t}-x(S_{i,t},A_{i,t})^{\top}\theta^{*}_{i,t}\right\}\big{]}$
$\leq\sum_{k=1}^{K}\frac{1}{k}\sum_{(i,t)\in\mathcal{O}_{k}\backslash\mathcal{O}_{k-1}}\big{[}$ $(\pi_{i,t}^{\star}-\pi_{i,t})x(S_{i,t},A_{i,t})^{\top}\theta^{*}_{i,t}\big{]}$
$+\sum_{k=1}^{K}\frac{1}{k}\sum_{(i,t)\in\mathcal{O}_{k}\backslash\mathcal{O}_{k-1}}\big{[}x(S_{i,t},a_{i,t}^{*})^{\top}\theta^{*}_{i,t}-x(S_{i,t},A_{i,t})^{\top}\theta^{*}_{i,t}\big{]}$
For the first term, $\sum_{k=1}^{K}\frac{1}{k}\sum_{(i,t)\in\mathcal{O}_{k}\backslash\mathcal{O}_{k-1}}\big{[}(\pi_{i,t}^{\star}-\pi_{i,t})x(S_{i,t},A_{i,t})^{\top}\theta^{*}_{i,t}\big{]}$, one can apply a nearly identical proof to that of Greenewald et al. (2017). For the second term we need a novel strategy, and thus focus on that term. The proof follows closely from Abeille & Lazaric (2017) with several adjustments. Assumption 6 implies that we only need to consider the unit ball $\mathcal{X}=\{\|x\|\leq 1\}$. Then the second term of the regret can be decomposed into
$\underbrace{\sum_{k=1}^{K}\frac{1}{k}\sum_{(i,t)\in\mathcal{O}_{k}\backslash\mathcal{O}_{k-1}}\left((\phi(x_{i,t}^{\star}))^{\top}\theta^{\star}-\phi(x_{i,t})^{\top}\tilde{\theta}_{k}\right)}_{R^{TS}(K)}+\underbrace{\sum_{k=1}^{K}\frac{1}{k}\sum_{(i,t)\in\mathcal{O}_{k}\backslash\mathcal{O}_{k-1}}\left(\phi(x_{i,t})^{\top}\tilde{\theta}_{k}-\phi(x_{i,t})^{\top}\theta^{\star}\right)}_{R^{RLS}(K)}$
where $\phi(x_{i,t}^{\star})$ is the context vector under the optimal action and $\theta^{\star}$ is the true parameter value. The first term is the regret due to the random deviations caused by sampling $\tilde{\theta}_{k}$ and whether it provides sufficient useful information about the true parameter $\theta^{\star}_{k}$. The second term is the concentration of the sampled term around the true linear model for the advantage function.
@@ F.3 Proof of Theorem 1
The first term describes the deviation of the TS linear predictor from the RLS one, while the second term describes the deviation of the RLS linear predictor from the true linear predictor. The first term is controlled by the construction of the sampling distribution $D^{TS}$, while the second term is controlled by the RLS estimate being a minimizer of the regularized cumulative squared error in (LABEL:eq:dml_finalform). In particular, the first term will be small when the TS estimate concentrates around the RLS one, while the second will be small when the RLS estimate concentrates around the true parameter vector. The next proposition gives a lower bound on the probability that, for all stages, both the RLS parameter vector concentrates around the true parameter vector and the TS parameter vector concentrates around the RLS one.
@@
$P(\hat{E}_{K})$ $=P(\cap_{k=1}^{K}\{\|\hat{\theta}_{k}-\theta^{\star}\|_{V_{k}}\leq\beta_{k}(\delta^{\prime})\})$
$=1-\sum_{k=1}^{K}P(\|\hat{\theta}_{k}-\theta^{\star}\|_{V_{k}}>\beta_{k}(\delta^{\prime}))$
@@
$+\sum_{k=1}^{K}\frac{1[E_{K}]}{k}\left[\sum_{(i,t)\in\mathcal{O}_{k}\backslash\mathcal{O}_{k-1}}|\phi(x_{i,t})^{\top}(\hat{\theta}_{k}-\theta^{*})|\right]$
@@
$\leq\left(\beta_{K}(\delta^{\prime})+\gamma_{K}(\delta^{\prime})\right)\sqrt{H_{K}}\sqrt{2c\left[2Kd\log\left(\gamma+\lambda M+3\frac{K+1}{16d}\right)-\log\det\left(V_{0}\right)\right]}$
@@
Leveraging Abeille & Lazaric (2017), Definition 1 lets us bound $R^{TS}(K)$ under the event $E_{k}$. Let $\phi(x_{i,t}^{\star})(\theta)=\arg\max_{x_{i,t}\in\cal X}\theta^{\top}\phi(x_{i,t})$. Then
@@
$R^{TS}(K)\leq\frac{4\gamma_{K}(\delta^{\prime})}{d}\left(\sqrt{\frac{8K}{\lambda}\log\left(\frac{4}{\delta}\right)}+\sqrt{4cH_{K}Kd\log\left(\gamma+\lambda M+3\frac{K+1}{16d}\right)-\log\det\left(V_{0}\right)}\right).$
@@
$\begin{split}&\left(\beta_{K}(\delta^{\prime})+\gamma_{K}(\delta^{\prime})\left[1+\frac{4}{d}\right]\right)\sqrt{4cH_{K}Kd\log\left(\gamma+\lambda M+3\frac{K+1}{16d}\right)-\log\det\left(V_{0}\right)}\\ +&\frac{4\gamma_{K}(\delta^{\prime})}{p}\sqrt{\frac{8K}{\lambda}\log\left(\frac{4}{\delta}\right)}\end{split}$
Appendix G Notation Guide
For convenience, we summarize below some key notation used in the main paper.
  • $a=0$: control action
  • $q$: number of non-baseline treatment arms
  • $i=1,2,\ldots$: index for individuals (later in the paper, we consider $N$ individuals)
  • $t=1,2,\ldots$: index for decision points (later in the paper, we consider $T$ decision points)
  • $S_{i,t}\in\mathcal{S}$: context vector observed for individual $i$ at decision point $t$
  • $A_{i,t}\in\{0,\ldots,K\}$: action chosen for individual $i$ at decision point $t$
  • $R_{i,t}\in\mathbb{R}$: reward observed for individual $i$ at decision point $t$
  • $r_{i,t}(s,a):=\mathbb{E}\left[R_{i,t}|S_{i,t}=s,A_{i,t}=a\right]$: conditional model for the observed reward given the state and context
  • $x(s,a)\in\mathbb{R}^{p\times 1}$: feature vector containing the state and action
  • $\theta_{i,t}\in\mathbb{R}^{p}$: vector of parameters that may depend on the individual $i$ and decision point $t$. Later $\theta_{i,t}$ is written as $\theta_{i,t}=\theta+\theta_{i}^{user}+\theta_{t}^{time}$, where $\theta_{i}^{user}$ is the individual-specific but time-invariant term and $\theta_{t}^{time}$ is a shared time-specific term
  • $\delta_{a>0}$: indicator function that takes the value $1$ if $a>0$ and 00 otherwise
  • $g_{t}(s)$: baseline reward function that is observed when individuals are randomized to receive no treatment
  • $\Delta_{i,t}(s,a):=r_{i,t}(s,a)-r_{i,t}(s,0)$: linear differential reward for any action $a>0$ and state $s$
  • $\mathcal{H}_{i,t}$: history up to decision point $t$ for individual $i$
  • $\pi_{i,t}(a|s)$: probability of action $a\in[K]$ given current context $s\in\mathcal{S}$ for a fixed (implicit) history
  • $\bar{a}_{i,t}\in[K]$: potential non-baseline arm that may be chosen if the baseline arm is not chosen
  • $f_{i,t}(s,a)$: working model for the true conditional mean $r_{i,t}(s,a)$
  • $\tilde{R}_{i,t}^{f}(s,\bar{a})$: pseudo-reward given state $S_{i,t}=s$ and potential arm $\bar{a}$, which has the same expectation as the differential reward, $\Delta_{i,t}(s_{i,t},\bar{a}_{i,t})$. This term is written as $\tilde{R}_{i,t}^{f}$ in some points in the main paper, with the state and action implied.
  • $\Delta_{i,t}^{f}(s,\bar{a})$: doubly robust estimator for the differential reward
  • $I_{m}(t)\subseteq\{1,...,t\}$: the $m$-th fold, of $M$ total folds, as assigned up to time $t$. $I^{\complement}_{m}(t)$ denotes its complement.
  • $\lambda$: parameter for ridge penalization, used in $\lambda||\theta||^{2}_{2}$
  • $\tilde{\sigma}^{2}_{i,t}=\pi_{i,t}(0|s_{i,t})\cdot(1-\pi_{i,t}(0|s_{i,t}))$: weights used in the penalized regression estimation, which are inversely proportional to $var(\tilde{R}^{f}_{i,t})$
  • $\left\{d(i,j):=||\theta_{i}-\theta_{j}||^{2}_{2}\right\}_{i\neq j}$: $L_{2}$-distances for network construction
  • $G=(V,E)$: graph with nodes $V$ and edges $E$. Each node corresponds to an individual, $V:=[N]$, and $(i,j)\in E$ for the smallest $M<<N$ distances.
  • B: incidence matrix where the element $B_{v,e}$ corresponds to the $v$-th vertex (individual) and $e$-th edge
  • $L$: Laplacian
  • $\mathcal{O}_{k}=\{(i,t):i\leq k\ \&\ t\leq k+1-i\}$: the set of observed time points across all individuals at stage $k$ of the sequential requirement setting
  • $G_{user}$: nearest neighbor graph characterizing proximity in the user domain
  • $G_{time}$: nearest neighbor graph characterizing proximity in the time domain item $\Theta_{k}=\textrm{vec}[\left(\theta,\theta_{1}^{\textrm{user}},\ldots,\theta_{k}^{\textrm{user}},\theta_{1}^{\textrm{time}},\ldots,\theta_{k}^{\textrm{time}}\right)]\in\mathbb{R}^{p(2k+1)}$: the set of all parameters, including the individual-specific time-invariant parameters and the shared time-specific parameters, at stage $k$ of the sequential recruitment setting
  • $M$, $D_{user}$, $D_{time}$: terms for bound on parameters and mean estimate.